indiaslayallday:

A number increases by 21 is equivalent to half the number increased by 18.5. Find the number.

Vocaloid:

@indiaslayallday hello

Vocaloid:

please let me know when you are ready to begin

Eiwoh2:

I know when: Never! ^^

AjStyles:

Answer is -5.

kittybasil:

No direct answers, thanks. Let's call the number "\(n\)" "A number increases by 21" \(n+21\) "is equivalent to" \(=\) "half the number" \(\frac{n}{2}\) "increased by 18.5" \(+18.5\) Therefore your equation is \[n+21=\frac{n}{2}+18.5\]


I'll be back with the solution, but if anyone wants to solve before that they can go ahead.

kittybasil:

Okay, back. Had lots of stuff to do.


\[n+21=\frac{n}{2}+18.5\] First let's get all the values with \(n\) on one side...\[n-0.5n+21=18.5\]Now we take all the non-variable numbers and put them on the other side.\[0.5n=18.5-21\]Solving this...\[0.5n=-2.5\]\[n=\frac{-2.5}{0.5}\]\[n=-5\]
\[n=-5\]

kittybasil:

And? Doesn't mean other people can't come in and look at the answer ☻